You have a $53,000 portfolio consisting of Intel, GE, and Con Edison. You put $21,200 in Intel, $13,200 in GE, and the rest in Con Edison. Intel, GE, and Con Edison have betas of 1.3, 1, and .8, respectively. What is your portfolio beta?

0.995

0.808

1.365

1.050
Answer #1

Investment in Con Edison=53000-(21200+13200)=$18600

Portfolio beta=Respective beta*Respective weight

=(21200/53000*1.3)+(13200/53000*1)+(18600/53000*0.8)

=1.050(Approx).